OpenStudy (anonymous):

Write the following rational numbers as ratios of integers. -5.0828282 and 7.421421421

OpenStudy (amistre64):

the .421 is 421/999

OpenStudy (anonymous):

Isn't there more to it or is that just the first step?

OpenStudy (amistre64):

1000(-5.08282...) = -5082.8282... - 10(-5.08282...) = -50.8282... ------------------------------ 1090=-5082-50 then divide of the left side

OpenStudy (amistre64):

990, forgot how to subtract

OpenStudy (amistre64):

this is a simple way to do it without having to resot to mathematical induction

OpenStudy (amistre64):

spose you get: .abcabcabc then there are 3 numbers repeating; put the repeaters over 3 9s abc/999

OpenStudy (anonymous):

wait where is 990 supposed to go? And what is final answer to -5.08282?

OpenStudy (amistre64):

rename: A=-5.08282... 1000 A = -5082.8282... - 10 A = -50.8282... ------------------------------ 990 A= -5082-50 A= -5132/990

OpenStudy (amistre64):

so A = A -5.08282.... = -5132/990

OpenStudy (amistre64):

reduce as needed

OpenStudy (anonymous):

alright i will reduce and thanks... what about the other problem or was that the final answer at the top?

OpenStudy (amistre64):

the final answer up top; just dont forget to add back in the 7 7+ 421/999 and reduce ;)

OpenStudy (anonymous):

wait so reduce and add 7 to the numerator or denominator or both haha

OpenStudy (amistre64):

its just a 7, nothing special about it. 7 + fraction

OpenStudy (amistre64):

and i need to recheck my calculations on the other one prolly

OpenStudy (anonymous):

alright haha can i put it in the calculator and also how do i show my work for this problem?

OpenStudy (amistre64):

depends on what level your spose to know; if this is for mathematical induction proofs, then youll have to go another route, otherwise id just do as i did

OpenStudy (amistre64):

the -5 is throwing me off i think; lets avoid it and go for the deci instead A = .08282 1000A = 82.8282 - 10A = .8282 ---------------- 990A = 82 A = 82/990 let me see what the wolf makes of that

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=-5-82%2F990 much better

OpenStudy (amistre64):

-(5+frac) so include back in the number part

OpenStudy (anonymous):

wait now i am confused more

OpenStudy (amistre64):

tell me where you got lost :)

OpenStudy (anonymous):

on both answers haha i thought i had the 1st one right then you said you might have got it wrong.

OpenStudy (amistre64):

the first one was good; it was the last one that the negative part was messing with me. 7 + frac = 7 + 421/999 -(5 + frac) = -(5 + 82/990 )

OpenStudy (anonymous):

alright so for the first one i got 7414/999 cause the fraction is added by 7 is that correct?

OpenStudy (amistre64):

yep that ones good

OpenStudy (anonymous):

Ok so the answer is 7141/999 and that can't be reduced... i am just making sure.

OpenStudy (anonymous):

sorry meant 7414/999